ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

9. Ordinary Differential Equations and Dynamical Systems

1991 MS Classification: 28, 58 Bezuglyi, Sergey, Kharkov, Ukraine and Kwiatkowski, Jan, Torun, Poland Full groups of minimal homeomorphisms of Cantor sets Let X be a Cantor set and let T be a minimal homeomorphism of X. The full group [T] generated by T is defined as the set {S ∈ Homeo(X) : S(x) = T^{n(x)}(x), x ∈ X}. We suggest a method of description of all homeomorphisms from [T] for a class of Cantor minimal systems (X, T). This method can be applied, for example, for odometers and Toeplitz flows.